Graph the following rational functions, give any equations for vertical,horizontal, or oblique asymptotes.
-6x-24+3x^2
------------
5-26x+ 5x^2

Answer by stanbon(48516)

=(6x-24+3x^2)/(5-26x+ 5x^2)
=(3^2+6x-24)/(5x^2-26x+5)
=3(x^2+2x-12)/(5x^2-25x-x+5)
=3(x^2+2x-12)/(5x(x-5)-(x-5))
=3(x^2+2x-12)/(x-5)(5x-1)
=========
This site will not support graphing your rational function.
Vertical asymptotes at x=5 and at x=1/5
Horizontal asymptote at y=3/5
-------------
Cheers,
Stan H.